OpenStudy (anonymous):

Create your own polynomial with a degree of 3 that when divided by x+2 has a remainder of -4.

OpenStudy (anonymous):

This is madness bears cant do math!

OpenStudy (dumbcow):

use remainder thm .... f(c) = remainder when divided by (x-c) so find a cubic function such that f(-2) = -4 f(x) = (x-a)(x-b)(x-c) try to get numbers a,b,c

OpenStudy (anonymous):

i made one, can you tell me if it's correct

OpenStudy (anonymous):

(6x^4-4x^3+5x^2-48x-16) / (x+2)

OpenStudy (dumbcow):

woah, its lot simpler just to leave it in factored form also that is a 4th degree ....not a cubic

OpenStudy (anonymous):

ohh okay

OpenStudy (dumbcow):

another tip: f(-2) = (-2-a)(-2-b)(-2-c) = -4 make the factors: 1,1,-4 since 1*1*-4 = -4 -2-a = 1 -2-b = 1 -2-c = -4